OpenStudy (hockeychick23):

Suppose a normal distribution has a mean of 98 and a standard deviation of 6. What is P(x<= 104) ?

jimthompson5910 (jim_thompson5910):

if the raw score is x = 104, then what is the z-score?

OpenStudy (hockeychick23):

ok then it would be 104-98=6/6=1

jimthompson5910 (jim_thompson5910):

That means P(X <= 104) is the same as P(Z <= 1)

jimthompson5910 (jim_thompson5910):

you would then use a table like this one http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf to find the area under the curve to the left of z = +1.00

jimthompson5910 (jim_thompson5910):

Tell me what you get.

OpenStudy (hockeychick23):

.8413

jimthompson5910 (jim_thompson5910):

very good, so because \[\Large P(Z \le 1) = 0.8413\] this means \[\Large P(X \le 104) = 0.8413\]

jimthompson5910 (jim_thompson5910):

well both approximate, but you get the idea

OpenStudy (hockeychick23):

ok thanks so much!!

jimthompson5910 (jim_thompson5910):

you're welcome